There were 662 tickets sold for a basketball game. Carholder's ticket is $1.75 and noncardholder's ticket is $2.50. The total money collected was $1399.25. How many of each kind of ticket was sold?
Cardholder's ticket=
Non-Cardholder's ticket=

Answer by htmentor(1343)

There were 662 tickets sold for a basketball game. Carholder's ticket is $1.75 and noncardholder's ticket is $2.50. The total money collected was $1399.25. How many of each kind of ticket was sold?
===========================================
Let c = the number of cardholder tickets
Then 662 - c = the number of non-cardholder tickets
The total amount collected is:
1.75c + 2.5(662-c) = 1399.25
Simplify and solve:
1.75c + 1655 - 2.5c = 1399.25
-0.75c = -255.75
c = 341
So the numbers are:
Cardholders = 341
Non-cardholders = 662 - 341 = 321